Consider a competitive market for pork with the quantity demanded (per year) at various prices are given as follows:

Price (dollars/kg)

Demand (million kg)

60

22

80

20

100

18

120

16

Calculate the price elasticity of demand when the price is $80/kg.  

Calculate the price elasticity of demand when the price is $00/kg.

Calculate the arc elasticity between price of $80/kg and $100/kg.
Answer #1

Ans 1)

Price elasticity=[(Q2-Q1)/(Q1)[(P1/(P2-P1)]
=[(20-22)/22][60/20]
=(-2/22)*(3)
=-6/22
=-0.27

Ans 2)

Price Elasticity of demand when Price=$80/Kg

Price elasticity=[(Q2-Q1)/(Q1)[(P1/(P2-P1)]
=[(18-20)/20][80/20]
=(-2/20)*(4)
=-8/20
=-0.4


Answer 3)

Price Elasticity of demand when Price=$80/Kg

Price elasticity=[(Q2-Q1)/(Q2+Q1)/2][(P1+P2)/2/(P2-P1)]
=[(18-20)/19][90/20]
=(-2/19)*(4.5)
=-9/19
=-0.473